Jefferson Lab Weekly Briefs August 10, 2011
Jefferson Lab Weekly Briefs
August 10, 2011
12 GeV Upgrade
With 15 weeks remaining in the Six Month Shutdown, 12 GeV installation work is progressing as planned. The first C-100 cryomodule is installed in the south linac of the accelerator, and all installation activities to support locking up the southern section of the machine in preparation for commissioning of the cryomodule are completing this week. Integrated 12 GeV logistical planning efforts, including material handling/transportation, space, storage and management of off-site facilities, are a continuous process that also incorporates fabrication and testing of components in preparation for installation during the next down period.
Hall A is continuing its installation of the g2p experiment. The septum magnet has been installed and aligned. Several new beamline magnets have been installed and are being cabled to their power supplies. The detector system for the High Resolution Spectrometer is being modified and prepared for the experiment.
In the injector, work on the injector experiment (PEPPo) continues, with fiber array detector checkout with the source and counter coincidence and data acquisition debugging. The 5 MeV dipole has been installed. In the east arc, vacuum hookup continues, and the polarity of the refurbished magnets in arcs 5, 7 and 9 has checked out as expected. In the west arc, alignment surveys continue throughout, with magnet stacks 13-15 recently installed and ready for survey; additionally, the 4-meter dipoles MXPAA 18-20 have been installed. In the south linac, warm region girders 2L21, 2L23 and 2L27 were installed; C100-1 was installed at 2L24 with the waveguide and beamline ion pumps connected to the new ion pump power supplies; the low-conductivity water system manifold brackets for 2L22, 2L23 and 2L24 were completed; the hose installs for 2L25 was completed; and the 2L25 high-power amplifier pedestal grounds were installed.
The FEL team reached its goal of finishing the gun high-voltage processing, although the group decided to limit operations at a bit lower voltage than originally planned. This week, the FEL team will cool down the cryounit to 2 Kelvin and reassemble the injector beamline.
Center for Theoretical and Computational Physics
In a new Theory Center paper, the pion mass and decay constant are calculated in lattice QCD for a range of quark masses. These results are analyzed using chiral perturbation theory (ChPT), allowing precise extractions of two important parameters in the ChPT Lagrangian. These coefficients appear in many phenomenological contexts, and the results obtained are comparable to extractions from experiment.
The electrical group completed the 6 GeV power supply re-installation and hook-up for the north and south access power supply rooms. This work was required due to infrastructure changes needed for the 12 GeV Upgrade. The east arc magnet interlock and electrical quality assurance checks were performed as magnets were re-installed in the tunnel. Polarity checks on the magnets for the south ends of both the east and west arcs is complete, and magnet installation continues on the north ends.
The electrical group also completed the CAMAC to VME conversion in the west arc service buildings W2 and W4. Network controls installation has now been completed in the south linac for radiofrequency zones SL23 and SL25. Zone SL24 controls, completed earlier, are being used to commission the new RF system in that zone. All eight klystrons in SL24 have been run up to approximately 10 kilowatts, exercising the new low-level RF controls. Detailed system checkout continues in SL24, while assembly and installation continues in SL23 and SL25. Waveguide installation preparation is underway in the tunnel to be ready for connection to the first C100 cryomodule this week. Work continues on the south linac Personnel Safety System upgrades needed for the new RF zones. Commissioning and certification are also scheduled this week.
The alignment group reports that they have completed the final survey for the east arc, south of the mid-arc gate. Additionally, the surveys for hook-up are continuing in the west arc and arc 10. A warm girder and cryomodule were aligned in the south linac last week.
Environment, Safety, Health and Quality
JLab requires that only safety equipment meeting the appropriate ANSI or NIOSH requirements is used. The ANSI marking relates to the American National Standards Institute, which organizes performance standards that safety equipment must meet. ANSI standards, signified by a stamp containing a Z plus a number, apply to the test protocols and requirements used for safety equipment, including shoes, toecaps, safety glasses, hard hats and hearing protection. Respiratory protection devices are tested by the National Institute of Occupational Safety and Health (NIOSH).
Using safety equipment that meets the appropriate ANSI or NIOSH standard ensures that JLab employees will receive the best protection available. Safety items that may be purchased through the Stockroom have been specified or accepted by Industrial Hygiene or Occupational Medicine. Safety equipment that is needed for a job must be obtained through the Stockroom or Industrial Hygiene. Buying safety items or non-ANSI (or similar-rated safety equipment) from catalogues or through the internet is not permitted. If you have questions or are unsure, contact the ESH&Q department for assistance.
JAG Luau Set for Wednesday, Aug. 17
Volunteers Needed for JAG Luau
United Way School Tools Drive
2011 Day of Caring Will 'Rock the Block'
Fed Travelers and Smartphone Apps
Engineering Offers Cryo Seminar
Tracking Thomas at JLab
Congratulations to Stephen Smith and Marie Ivanco, who were the first to correctly identify the Aug. 3 location. Honorable mentions go to Michael Spata, Percy Harrell, Lori Zukerman, Dana Cochran, Michael Haddox-Schatz, Karen Bartek, Lawrence Ferbee, Jason Willoughby, Kevin Jordan, DeAnn Maddox, Dave Fazenbaker, Gina Dixon, Narciso Gomez, Sharon Hay, Janet Prater, Jodi Patient, Dianne Napier, Evelyn Akers, Russell Mammei, Linda Ceraul, Luci Collins, John Kelly, Tonya Evans, Judy Hill, Cela Callaghan, Chandra Gilchrist, Connie Adams, Sebastien Procureur, Debra Brand, Casey Heck, Marie Powers, Mike Koch, Andre Newman, Mike Martin, Ron Bartek, Michelle Shinn, Andy Kowalski, Doug Higinbotham and Dick Owen. Check out the Tracking Thomas webpage for a better view of his last location and this week's new mystery photo.
JLab Calendar of Events